how do i find the answer for 2x^2-11x+12=0 i get stuck i know that you have to factor 12 to get -11 which be -12, 1 but from there i dont know what to do can you please help me?
Answer by mananth(16946)

2x^2-11x+12=0
..
2*12 = 24
a+b = -11
ab=24
-8,-3
...
2x^2-8x-3+12
2x(x-4)-3(x-4)=0
(x-4)(2x-3)=0
x= 4 OR x 3/2
...
m.ananth@hotmail.ca